OpenStudy (anonymous):

simplify 4^sqroot 80b^4c^3

OpenStudy (anonymous):

\[\sqrt{80b ^{4}c ^{3}}=\sqrt{4^{2}b ^{4}c*c ^{2}}=4b ^{2}c \sqrt{5c}\] so, the answer is: \[4^{4b ^{2}c \sqrt{5c}}\]